# Key Words.: binaries: spectroscopic – binaries: close – stars: early-type – stars: massive – Magellanic Clouds1 Introduction

### v1

@@ Acknowledgements.
Appendix A Tables
BLOeM ID | SpT | n | RV | Binarity | Period | EB | OGLE_P | Comments | |
km s-1 | [d] | [d] | |||||||
1-002 | B0 IV: | 9 | 66.33 | 31.21 | SB1 | 48.87 | |||
1-010 | B1.5 III: | 9 | 9.76 | 3.17 | RVvar | 1.48 | |||
1-013 | B1.5 III: | 9 | 69.94 | 12.29 | SB1 | 2.61 | |||
1-015 | B0.5 III | 9 | 74.14 | 31.09 | SB1 | 22.11 | |||
1-020 | B0 III | 9 | 93.23 | 41.72 | SB1 | 38.34 | |||
1-022 | B1 III | 9 | 8.52 | 1.70 | RVcst | 1.29 | |||
1-032 | B1.5 III: | 9 | 121.15 | 38.40 | SB1 | 3.05 | |||
1-035 | B1.5 III-II | 9 | 8.60 | 2.96 | RVvar | 2.04 | |||
1-037 | B1.5 III | 9 | 317.85 | 47.05 | SB2 | 4.99 | y | 3.324534 | Probable P |
1-046 | B1 III | 9 | 9.16 | 3.47 | RVvar | 4.61 | |||
1-047 | B1 III-II | 9 | 42.68 | 20.35 | SB1 | 203.12 | |||
1-048 | B2 III-II | 9 | 5.59 | 1.89 | RVcst | 2.09 | |||
1-055 | B1.5 III-II | 9 | 112.08 | 11.22 | SB2 | 3.04 | Potential Algol | ||
1-059 | B0 IV | 9 | 6.81 | 1.67 | RVcst | 6.55 | |||
1-067 | B0.7 III | 6 | 3.21 | 1.59 | RVcst | 1.55 | |||
1-069 | B0.7 III | 9 | 106.60 | 16.38 | SB2 | 2.08 | Misident. | ||
1-071 | B0.7 III | 9 | 9.14 | 3.44 | RVvar | 1.30 | |||
1-081 | B1.5: III: neb | 3 | 23.30 | 2.11 | RVvar | 467.41 | |||
1-082 | B1.5 III-II | 9 | 27.34 | 7.92 | SB1 | 7.77 | |||
1-083 | B0.5: V: | 9 | 190.41 | 15.75 | SB2 | 1.80 | Too blended - SB3? | ||
1-086 | B0 IV: | 9 | 56.70 | 19.89 | SB1 | 51.53 | |||
1-088 | B1.5 III | 9 | 131.43 | 13.53 | SB2 | 1.27 | Misident. | ||
1-101 | B1.5 III-II | 9 | 15.12 | 2.47 | RVvar | 4.62 | |||
1-107 | B0.5 III-II | 9 | 2.78 | 1.14 | RVcst | 4.26 | |||
2-006 | B1.5 III | 9 | 23.73 | 8.44 | SB1 | 1.62 | |||
2-010 | B1.5 III-II | 9 | 5.89 | 3.62 | RVvar | 1.51 | |||
2-023 | B2.5 III | 9 | 28.67 | 2.61 | SB2 | 4.67 | Potential Algol | ||
2-027 | B1.5 III neb | 8 | 34.93 | 5.25 | SB1 | 3.91 | |||
2-033 | B1 III-II | 9 | 31.03 | 7.86 | SB1 | 1.26 | |||
2-042 | B2 IV: | 9 | 29.66 | 8.33 | SB1 | 1.53 | |||
2-045 | B2 III: | 9 | 40.54 | 12.78 | SB1 | 7.32 | |||
2-049 | B1.5: III | 9 | 20.37 | 5.66 | SB1 | 500.00 | |||
2-051 | B2.5 IV: | 9 | 62.59 | 9.91 | SB1 | 2.00 | y | 2.336134 | |
2-052 | B0.5 V: | 9 | 204.57 | 40.43 | SB2 | 3.79 | y | 3.820998 | Potential Algol - SB3? |
2-053 | B2 III | 9 | 30.98 | 5.15 | SB1 | 3.45 | |||
2-062 | B1 III-II | 9 | 7.80 | 4.00 | RVvar | 1.29 | |||
2-064 | B1.5 V-III | 9 | 269.74 | 34.18 | SB2 | 2.58 | y | 2.57832 | SB3/LPV? |
2-094 | B0.5 III-II | 9 | 3.14 | 1.94 | RVcst | 1.51 | |||
2-095 | B0 IV | 9 | 2.24 | 1.11 | RVcst | 2.04 | |||
2-099 | B0 III: | 9 | 34.32 | 7.48 | SB1 | 3.42 | |||
2-100 | B0 V | 9 | 12.09 | 4.81 | RVvar | 1.24 | |||
2-103 | B0.5 V: | 9 | 135.14 | 11.34 | SB2 | 3.35 | Potential Algol | ||
2-107 | B1.5 III-II | 9 | 75.41 | 41.80 | SB1 | 4.44 | |||
3-003 | B1.5 III | 9 | 9.15 | 1.93 | RVcst | 1.66 | |||
3-005 | B1 III | 4 | 2.00 | 0.64 | RVcst | 40.37 | |||
3-015 | B2 III | 9 | 17.13 | 3.87 | RVvar | 6.75 | |||
3-020 | B0.5 III | 9 | 22.85 | 1.49 | SB2 | 3.36 | Potential Algol | ||
3-026 | B1.5 II-III | 9 | 28.23 | 12.12 | SB1 | 500.00 | |||
3-034 | B0 III | 9 | 32.70 | 11.37 | SB1 | 56.97 | |||
3-036 | B0.7 III-II | 9 | 3.80 | 2.42 | RVvar | 1.25 | |||
3-048 | B0.7 III | 9 | 6.02 | 2.60 | RVvar | 1.26 | |||
3-050 | B1.5 III | 9 | 75.69 | 27.41 | SB1 | 4.41 | y | 4.424906 | |
3-055 | B0.7 IV | 9 | 3.38 | 1.20 | RVcst | 3.63 | |||
3-067 | B0.2 IV | 9 | 274.24 | 28.35 | SB2 | 2.01 | y | 2.000392 | Misident.? |
3-071 | B0.2 V | 9 | 63.14 | 14.30 | SB1 | 153.21 | |||
3-072 | B1 III | 9 | 157.60 | 28.01 | SB2 | 3.06 | y | 2.489125 | Misident. - SB3? |
3-079 | B1.5 III | 9 | 196.18 | 47.99 | SB2 | 2.95 | Misident. - Good P | ||
3-080 | B1 III-II | 9 | 52.93 | 24.84 | SB1 | 163.70 | |||
3-086 | B1 III: | 9 | 40.56 | 10.12 | SB2 | 8.38 | Potential Algol - Possible SB3 | ||
3-089 | B2 III: | 9 | 97.92 | 20.92 | SB1 | 21.75 | |||
3-091 | B1 III-II | 9 | 5.84 | 2.25 | RVvar | 1.39 | |||
3-105 | B0.5 III-II | 9 | 2.03 | 1.31 | RVcst | 2.00 | |||
3-108 | B0 III | 9 | 1.89 | 0.82 | RVcst | 8.81 | |||
4-005 | B1 III-II | 9 | 122.27 | 20.48 | SB2 | 1.37 | Misident. - Widths overestimated | ||
4-007 | B1.5 III | 9 | 8.66 | 4.09 | RVvar | 1.48 | |||
4-009 | B2 III | 9 | 23.21 | 4.02 | SB2 | 1.42 | y | 1.417171 | Possible SB3 |
4-012 | B2 III | 9 | 12.13 | 3.02 | RVvar | 1.92 | |||
4-013 | B0.2 IV | 9 | 7.08 | 2.11 | RVvar | 4.55 | |||
4-014 | B0 III | 9 | 9.08 | 4.38 | RVvar | 1.34 | |||
4-016 | B0.7 III | 9 | 28.69 | 6.17 | SB1 | 1.66 | |||
4-025 | B1 III | 9 | 9.52 | 2.87 | RVvar | 3.34 | |||
4-029 | B1.5 IV | 9 | 54.52 | 12.14 | SB1 | 1.53 | y | 1.522977 | |
4-031 | B0 V neb | 8 | 68.62 | 5.68 | SB1 | 1.93 | |||
4-033 | B0.7 III | 9 | 7.27 | 2.08 | RVvar | 1.27 | |||
4-038 | B1.5 III-II | 9 | 4.85 | 2.24 | RVvar | 1.62 | |||
4-044 | B1.5 III neb | 8 | 11.93 | 2.72 | RVvar | 2.08 | |||
4-052 | B1.5 III: | 9 | 5.81 | 1.86 | RVcst | 1.84 | |||
4-053 | B0 V neb | 8 | 53.27 | 15.29 | SB1 | 144.95 | |||
4-056 | B0 V: neb | 9 | 41.33 | 3.57 | SB2 | 41.50 | y | 41.494224 | Bstr candidate - Possible SB3 - neb? |
4-059 | B1: III: | 9 | 128.06 | 12.41 | SB2 | 50.55 | y | 1.94869 | Misident. |
4-062 | B2: III: | 9 | 183.31 | 15.38 | SB2 | 2.68 | Potential Algol | ||
4-064 | B2 IV: | 9 | 14.51 | 3.13 | RVvar | 63.04 | |||
4-081 | B0.5: V | 9 | 149.19 | 12.44 | SB2 | 1.48 | Misident. | ||
4-088 | B0.5: IV: | 9 | 22.24 | 3.50 | RVvar | 5.18 | |||
4-092 | B2 III | 9 | 69.52 | 8.07 | SB2 | 1.49 | Misident. - Too blended | ||
4-093 | B1 III: | 9 | 11.74 | 6.59 | RVvar | 2.64 | |||
4-094 | B1 III | 9 | 31.50 | 12.55 | SB1 | 18.40 | |||
4-095 | B2 III: | 9 | 245.78 | 29.90 | SB2 | 3.41 | y | 1.573979 | P=2.8? - SB3? |
4-097 | B2 III: | 9 | 128.37 | 45.05 | SB1 | 3.19 | |||
4-099 | B0.7 III | 9 | 7.34 | 2.38 | RVvar | 1.92 | |||
4-100 | B1 III | 9 | 13.63 | 4.72 | RVvar | 7.13 | |||
4-106 | B2-2.5 III | 9 | 75.23 | 28.16 | SB1 | 4.60 | |||
4-108 | B1.5 III: | 5 | 5.31 | 1.32 | RVcst | 4.85 | |||
4-109 | B0.7 III | 9 | 5.30 | 1.57 | RVcst | 1.27 | |||
4-115 | B2 III: | 9 | 26.92 | 4.45 | SB1 | 2.07 | |||
4-116 | B1.5 III: | 9 | 33.62 | 5.76 | SB1 | 34.36 | |||
5-002 | B2 IV | 9 | 184.80 | 10.02 | SB2 | 2.06 | y | 0.91773 | Good P |
5-003 | B2.5 III | 9 | 19.84 | 3.82 | RVvar | 10.28 | |||
5-004 | B2.5 III-II | 9 | 12.76 | 3.09 | RVvar | 1.25 | |||
5-011 | B2 III: | 9 | 22.74 | 3.84 | RVvar | 2.84 | |||
5-013 | B2 III | 9 | 254.06 | 28.15 | SB2 | 1.56 | y | 3.509807 | Misident. |
5-015 | B2 III: | 9 | 94.84 | 20.87 | SB1 | 500.00 | |||
5-017 | B0.2 IV | 9 | 11.27 | 2.05 | RVvar | 1.40 | |||
5-019 | B2 III | 9 | 44.77 | 20.97 | SB1 | 35.45 | |||
5-024 | B0.7 III | 9 | 1.66 | 1.37 | RVcst | 16.34 | |||
5-025 | B1 III | 5 | 8.19 | 1.72 | RVcst | 1.49 | |||
5-027 | B1.5 III: | 9 | 167.13 | 25.61 | SB2 | 1.47 | Good P | ||
5-028 | B0 V | 9 | 27.40 | 3.53 | RVvar | 72.63 | |||
5-030 | B1.5 III: | 9 | 39.21 | 5.84 | SB2 | 1.32 | y | 2.638655 | Misident. |
5-033 | B1 III | 9 | 4.97 | 1.54 | RVcst | 2.43 | |||
5-037 | B0.5 V: | 9 | 53.73 | 12.87 | SB2 | 3.80 | y | 2.722014 | Misident. |
5-038 | B1.5 III-II | 9 | 19.79 | 11.81 | RVvar | 6.09 | |||
5-039 | B1.5 III | 9 | 101.10 | 32.26 | SB1 | 33.35 | |||
5-049 | B0.2 IV | 9 | 66.64 | 10.56 | SB1 | 4.98 | y | 9.935441 | |
5-051 | B0.2 III | 9 | 90.21 | 4.88 | SB1 | 9.49 | |||
5-057 | B1 III: | 9 | 42.82 | 4.83 | SB2 | 14.21 | Possible SB3 - Similar to 3-086 | ||
5-058 | B1 III-II | 9 | 2.00 | 0.82 | RVcst | 1.32 | |||
5-059 | B0.7 IV | 9 | 8.24 | 2.08 | RVvar | 8.70 | |||
5-061 | B2 III: | 9 | 14.61 | 2.81 | RVvar | 500.00 | |||
5-062 | B1.5 | 9 | 296.70 | 49.37 | SB3 | 79.96 | y | 5.24094 | Affected by 3rd light |
5-063 | B0 V: | 4 | 264.01 | 57.31 | SB2 | 1.84 | Too few epochs | ||
5-065 | B0 IV | 9 | 15.08 | 9.04 | RVvar | 500.00 | |||
5-068 | B1 III-II | 9 | 17.27 | 8.18 | RVvar | 1.22 | |||
5-070 | B2 III: | 9 | 9.10 | 2.41 | RVvar | 1.34 | |||
5-074 | B0.7 IV | 9 | 22.11 | 8.43 | SB1 | 1.65 | |||
5-076 | B1.5 III: | 9 | 22.83 | 3.34 | RVvar | 10.17 | |||
5-078 | B1.5 III: | 9 | 35.22 | 10.61 | SB1 | 1.33 | |||
5-079 | B2 III: | 9 | 52.02 | 16.37 | SB1 | 24.06 | |||
5-080 | B2 III: | 9 | 322.04 | 42.61 | SB2 | 2.51 | Misident. | ||
5-081 | B0.7 III | 9 | 9.47 | 1.46 | RVcst | 1.27 | |||
5-082 | B1.5 III | 9 | 14.69 | 2.13 | RVvar | 1.45 | |||
5-085 | B0.7 III | 9 | 6.28 | 2.26 | RVvar | 1.26 | |||
5-088 | B1.5 III-II | 9 | 5.46 | 1.69 | RVcst | 1.58 | |||
5-095 | B1.5 III: | 9 | 66.24 | 18.55 | SB1 | 102.95 | |||
5-096 | B1.5 III-II | 9 | 6.03 | 2.08 | RVvar | 1.22 | |||
5-100 | B0.2 V | 9 | 116.49 | 23.18 | SB2 | 5.61 | y | 2.044868 | Misident. |
5-110 | B1 III | 9 | 154.20 | 6.35 | SB2 | 1.26 | Unconstrained | ||
[…]

### v4

@@ Acknowledgements.
Appendix A Additional figures
A.1 Spatial distribution of the B III/V sample
![[Uncaptioned image]](x10.png)
A.2 Hertzsprung-Russell diagram
Figure 20 shows the Hertzsprung-Russell diagram (HRD) of the full BLOeM sample, colour-coded by their different spectral type regimes. Each of the subsamples has been investigated independently: the O-type stars by Sana et al. (2024), the B-type dwarfs and giants in this work, the early B-type supergiants (BSGs) by Britavskiy et al. (2025), the Oe/Be stars by Bodensteiner et al. (2025), and the late BAF supergiants by Patrick et al. (2025). All these works have investigated the multiplicity fraction of their respective samples, and there is a clear trend with evolutionary phase. The two samples with the less evolved content, non-supergiant O- and B-type stars, have the highest multiplicity fractions, with found by Sana et al. (2024) and found in this work. The fraction rapidly decreases for early BSGs, reaching . Furthermore, Britavskiy et al. (2025) reports a drop in the observed binary fraction at spectral types later than B2. Looking at the HRD of Fig. 20, we can see that the cooler BSGs have values above 4.5 dex. Therefore, it is likely that the drop in the binary fraction is related to the evolutionary phase of these stars; namely, the large radii of the BSGs have induced binary interactions, effectively removing the close binaries. This is supported by the study of Bodensteiner et al. (2025), who found a binary fraction of 18-32% among the emission-line stars. The stark contrast in the multiplicity properties of Oe/Be stars and OB-type stars strongly suggests that the former are mostly binary interaction products, as proposed by Bodensteiner et al. (2025). In many cases, the initially more massive star has lost a large fraction of its envelope mass and produces RV variations on the now more massive companion that are too small to detect, leading to a lower observed binary fraction. The more evolved stars in the BLOeM sample, the BAF supergiants, present the lowest binary fraction of all, with less than of the late BSGs and of the A- and F-type stars showing binary signatures (Patrick et al. 2025). Although the intrinsic binary fractions for the supergiant samples are highly uncertain due to evolutionary aspects and the unclear role of intrinsic variability, the trend is clear. Binary evolution strongly impacts the evolutions of massive stars, as is evidenced by the large drop in the binary fraction from the MS to the late supergiant phases; if the intrinsic binary fraction of AF supergiants is really below 20%, then close to 60% of massive stars must have gone through envelope stripping or mergers before reaching the AF supergiant phase, and now appear as single stars.
Appendix B Radial velocity uncertainties
Appendix C Notes on individual systems
C.1 Confirmed SB3 systems
BLOeM 5-062:
(OGLE-SMC-ECL-5096) Characterised by a narrow absorption feature moving with a large RV. As shown in Fig. 23, the core of the line is not well fitted with a Gaussian profile, and in epochs 5 and 6 there are two other absorption features at both sides of the narrow absorption. Our inspection of the photometric data shows that this is likely a doubly eclipsing system, with a potential outer period of 77.2 d, close to the spectroscopic period (80 d). The OGLE (inner) photometric period is 5.2 d.
BLOeM 6-062:
BLOeM 7-057:
C.2 Other eclipsing binaries
BLOeM 1-037:
BLOeM 2-051:
BLOeM 3-050:
BLOeM 3-072:
BLOeM 4-095:
BLOeM 5-002:
BLOeM 5-030:
BLOeM 6-115:
BLOeM 7-032:
BLOeM 7-060:
BLOeM 8-029:
BLOeM 8-105:
Appendix D Tables
BLOeM ID | SpT | n | RV | Binarity | Period | P_sig | EB | OGLE_P | Comments | |
km s-1 | [d] | [d] | ||||||||
BLOeM_1-002 | B0 IV: | 9 | 66.33 | 31.21 | SB1 | 48.87 | T | – | – | – |
BLOeM_1-010 | B1.5 III: | 9 | 9.76 | 3.17 | RVvar | 1.48 | NS | – | – | – |
BLOeM_1-013 | B1.5 III: | 9 | 69.94 | 12.29 | SB1 | 2.61 | T | – | – | – |
BLOeM_1-015 | B0.5 III | 9 | 74.14 | 31.09 | SB1 | 22.11 | NS | – | – | – |
BLOeM_1-020 | B0 III | 9 | 93.23 | 41.72 | SB1 | 38.34 | T | – | – | – |
BLOeM_1-022 | B1 III | 9 | 8.52 | 1.70 | RVcst | 1.29 | NS | – | – | – |
BLOeM_1-032 | B1.5 III: | 9 | 121.15 | 38.40 | SB1 | 3.05 | S | – | – | – |
BLOeM_1-035 | B1.5 III-II | 9 | 8.60 | 2.96 | RVvar | 2.04 | NS | – | – | – |
BLOeM_1-037 | B1.5 III | 9 | 317.85 | 47.05 | SB2 | 4.99 | NS | y | 3.324534 | – |
BLOeM_1-046 | B1 III | 9 | 9.16 | 3.47 | RVvar | 4.61 | NS | – | – | – |
BLOeM_1-047 | B1 III-II | 9 | 42.68 | 20.35 | SB1 | 203.12 | S | – | – | – |
BLOeM_1-048 | B2 III-II | 9 | 5.59 | 1.89 | RVcst | 2.09 | NS | – | – | – |
BLOeM_1-055 | B1.5 III-II | 9 | 112.08 | 11.22 | SB2 | 3.04 | S | – | – | Potential Algol |
BLOeM_1-059 | B0 IV | 9 | 6.81 | 1.67 | RVcst | 6.55 | NS | – | – | – |
BLOeM_1-067 | B0.7 III | 6 | 3.21 | 1.59 | RVcst | 1.55 | T | – | – | – |
BLOeM_1-069 | B0.7 III | 9 | 106.60 | 16.38 | SB2 | 2.08 | NS | – | – | Misident. |
BLOeM_1-071 | B0.7 III | 9 | 9.14 | 3.44 | RVvar | 1.30 | NS | – | – | – |
BLOeM_1-081 | B1.5: III: neb | 3 | 23.30 | 2.11 | RVvar | 467.41 | NS | – | – | – |
BLOeM_1-082 | B1.5 III-II | 9 | 27.34 | 7.92 | SB1 | 7.77 | NS | – | – | – |
BLOeM_1-083 | B0.5: V: | 9 | 190.41 | 15.75 | SB2 | 1.80 | NS | – | – | Too blended - SB3? |
BLOeM_1-086 | B0 IV: | 9 | 56.70 | 19.89 | SB1 | 51.53 | NS | – | – | – |
BLOeM_1-088 | B1.5 III | 9 | 131.43 | 13.53 | SB2 | 1.27 | NS | – | – | Misident. |
BLOeM_1-101 | B1.5 III-II | 9 | 15.12 | 2.47 | RVvar | 4.62 | NS | – | – | – |
BLOeM_1-107 | B0.5 III-II | 9 | 2.78 | 1.14 | RVcst | 4.26 | NS | – | – | – |
BLOeM_2-006 | B1.5 III | 9 | 23.73 | 8.44 | SB1 | 1.62 | NS | – | – | – |
BLOeM_2-010 | B1.5 III-II | 9 | 5.89 | 3.62 | RVvar | 1.51 | NS | – | – | – |
BLOeM_2-023 | B2.5 III | 9 | 28.67 | 2.61 | SB2 | 4.67 | S | – | – | Potential Algol |
BLOeM_2-027 | B1.5 III neb | 8 | 34.93 | 5.25 | SB1 | 3.91 | NS | – | – | – |
BLOeM_2-033 | B1 III-II | 9 | 31.03 | 7.86 | SB1 | 1.26 | S | – | – | – |
BLOeM_2-042 | B2 IV: | 9 | 29.66 | 8.33 | SB1 | 1.53 | NS | – | – | – |
BLOeM_2-045 | B2 III: | 9 | 40.54 | 12.78 | SB1 | 7.32 | S | – | – | – |
BLOeM_2-049 | B1.5: III | 9 | 20.37 | 5.66 | SB1 | 500.00 | NS | – | – | – |
BLOeM_2-051 | B2.5 IV: | 9 | 62.59 | 9.91 | SB1 | 2.00 | S | y | 2.336134 | SB2? |
BLOeM_2-052 | B0.5 V: | 9 | 204.57 | 40.43 | SB2 | 3.79 | T | y | 3.820998 | Potential Algol - SB3? |
BLOeM_2-053 | B2 III | 9 | 30.98 | 5.15 | SB1 | 3.45 | NS | – | – | – |
BLOeM_2-062 | B1 III-II | 9 | 7.80 | 4.00 | RVvar | 1.29 | NS | – | – | – |
BLOeM_2-064 | B1.5 V-III | 9 | 269.74 | 34.18 | SB2 | 2.58 | S | y | 2.57832 | SB3/LPV? |
BLOeM_2-094 | B0.5 III-II | 9 | 3.14 | 1.94 | RVcst | 1.51 | NS | – | – | – |
BLOeM_2-095 | B0 IV | 9 | 2.24 | 1.11 | RVcst | 2.04 | NS | – | – | – |
BLOeM_2-099 | B0 III: | 9 | 34.32 | 7.48 | SB1 | 3.42 | NS | – | – | – |
BLOeM_2-100 | B0 V | 9 | 12.09 | 4.81 | RVvar | 1.24 | NS | – | – | – |
BLOeM_2-103 | B0.5 V: | 9 | 135.14 | 11.34 | SB2 | 3.35 | S | – | – | Potential Algol |
BLOeM_2-107 | B1.5 III-II | 9 | 75.41 | 41.80 | SB1 | 4.44 | S | – | – | – |
BLOeM_3-003 | B1.5 III | 9 | 9.15 | 1.93 | RVcst | 1.66 | NS | – | – | – |
BLOeM_3-005 | B1 III | 4 | 2.00 | 0.64 | RVcst | 40.37 | NS | – | – | – |
BLOeM_3-015 | B2 III | 9 | 17.13 | 3.87 | RVvar | 6.75 | NS | – | – | – |
BLOeM_3-020 | B0.5 III | 9 | 22.85 | 1.49 | SB2 | 3.36 | T | – | – | Potential Algol |
BLOeM_3-026 | B1.5 II-III | 9 | 28.23 | 12.12 | SB1 | 500.00 | NS | – | – | – |
BLOeM_3-034 | B0 III | 9 | 32.70 | 11.37 | SB1 | 56.97 | NS | – | – | – |
BLOeM_3-036 | B0.7 III-II | 9 | 3.80 | 2.42 | RVvar | 1.25 | T | – | – | – |
BLOeM_3-048 | B0.7 III | 9 | 6.02 | 2.60 | RVvar | 1.26 | NS | – | – | – |
BLOeM_3-050 | B1.5 III | 9 | 75.69 | 27.41 | SB1 | 4.41 | S | y | 4.424906 | – |
BLOeM_3-055 | B0.7 IV | 9 | 3.38 | 1.20 | RVcst | 3.63 | NS | – | – | – |
BLOeM_3-067 | B0.2 IV | 9 | 274.24 | 28.35 | SB2 | 2.01 | S | y | 2.000392 | Misident.? |
BLOeM_3-071 | B0.2 V | 9 | 63.14 | 14.30 | SB1 | 153.21 | NS | – | – | – |
BLOeM_3-072 | B1 III | 9 | 157.60 | 28.01 | SB2 | 3.06 | NS | y | 2.489125 | Misident. - SB3? |
BLOeM_3-079 | B1.5 III | 9 | 196.18 | 47.99 | SB2 | 2.95 | S | – | – | – |
BLOeM_3-080 | B1 III-II | 9 | 52.93 | 24.84 | SB1 | 163.70 | S | – | – | – |
BLOeM_3-086 | B1 III: | 9 | 40.56 | 10.12 | SB2 | 8.38 | T | – | – | Potential Algol - Possible SB3 |
BLOeM_3-089 | B2 III: | 9 | 97.92 | 20.92 | SB1 | 21.75 | NS | – | – | – |
BLOeM_3-091 | B1 III-II | 9 | 5.84 | 2.25 | RVvar | 1.39 | NS | – | – | – |
BLOeM_3-105 | B0.5 III-II | 9 | 2.03 | 1.31 | RVcst | 2.00 | NS | – | – | – |
BLOeM_3-108 | B0 III | 9 | 1.89 | 0.82 | RVcst | 8.81 | NS | – | – | – |
BLOeM_4-005 | B1 III-II | 9 | 122.27 | 20.48 | SB2 | 1.37 | NS | – | – | Narrow cores - Misident. |
BLOeM_4-007 | B1.5 III | 9 | 8.66 | 4.09 | RVvar | 1.48 | NS | – | – | – |
BLOeM_4-009 | B2 III | 9 | 23.21 | 4.02 | SB2 | 1.42 | NS | y | 1.417171 | Possible SB3 |
BLOeM_4-012 | B2 III | 9 | 12.13 | 3.02 | RVvar | 1.92 | NS | – | – | – |
BLOeM_4-013 | B0.2 IV | 9 | 7.08 | 2.11 | RVvar | 4.55 | NS | – | – | – |
BLOeM_4-014 | B0 III | 9 | 9.08 | 4.38 | RVvar | 1.34 | NS | – | – | – |
BLOeM_4-016 | B0.7 III | 9 | 28.69 | 6.17 | SB1 | 1.66 | T | – | – | – |
BLOeM_4-025 | B1 III | 9 | 9.52 | 2.87 | RVvar | 3.34 | NS | – | – | – |
BLOeM_4-029 | B1.5 IV | 9 | 54.52 | 12.14 | SB1 | 1.53 | S | y | 1.522977 | – |
BLOeM_4-031 | B0 V neb | 8 | 68.62 | 5.68 | SB1 | 1.93 | S | – | – | – |
BLOeM_4-033 | B0.7 III | 9 | 7.27 | 2.08 | RVvar | 1.27 | NS | – | – | – |
BLOeM_4-038 | B1.5 III-II | 9 | 4.85 | 2.24 | RVvar | 1.62 | NS | – | – | – |
BLOeM_4-044 | B1.5 III neb | 8 | 11.93 | 2.72 | RVvar | 2.08 | NS | – | – | – |
BLOeM_4-052 | B1.5 III: | 9 | 5.81 | 1.86 | RVcst | 1.84 | NS | – | – | – |
BLOeM_4-053 | B0 V neb | 8 | 53.27 | 15.29 | SB1 | 144.95 | NS | – | – | – |
BLOeM_4-056 | B0 V: neb | 9 | 41.33 | 3.57 | SB2 | 41.50 | T | y | 41.494224 | Bstr candidate - Possible SB3 |
BLOeM_4-059 | B1: III: | 9 | 128.06 | 12.41 | SB2 | 50.55 | NS | y | 1.94869 | Misident. |
BLOeM_4-062 | B2: III: | 9 | 183.31 | 15.38 | SB2 | 2.68 | S | – | – | Potential Algol |
BLOeM_4-064 | B2 IV: | 9 | 14.51 | 3.13 | RVvar | 63.04 | NS | – | – | – |
BLOeM_4-081 | B0.5: V | 9 | 149.19 | 12.44 | SB2 | 1.48 | NS | – | – | Misident. |
BLOeM_4-088 | B0.5: IV: | 9 | 22.24 | 3.50 | RVvar | 5.18 | NS | – | – | – |
BLOeM_4-092 | B2 III | 9 | 69.52 | 8.07 | SB2 | 1.49 | NS | – | – | Misident. - Too blended |
BLOeM_4-093 | B1 III: | 9 | 11.74 | 6.59 | RVvar | 2.64 | NS | – | – | – |
BLOeM_4-094 | B1 III | 9 | 31.50 | 12.55 | SB1 | 18.40 | NS | – | – | – |
BLOeM_4-095 | B2 III: | 9 | 245.78 | 29.90 | SB2 | 3.41 | NS | y | 1.573979 | P=2.8 d? - SB3? |
BLOeM_4-097 | B2 III: | 9 | 128.37 | 45.05 | SB1 | 3.19 | S | – | – | – |
BLOeM_4-099 | B0.7 III | 9 | 7.34 | 2.38 | RVvar | 1.92 | S | – | – | – |
BLOeM_4-100 | B1 III | 9 | 13.63 | 4.72 | RVvar | 7.13 | NS | – | – | – |
BLOeM_4-106 | B2-2.5 III | 9 | 75.23 | 28.16 | SB1 | 4.60 | S | – | – | – |
BLOeM_4-108 | B1.5 III: | 5 | 5.31 | 1.32 | RVcst | 4.85 | NS | – | – | – |
BLOeM_4-109 | B0.7 III | 9 | 5.30 | 1.57 | RVcst | 1.27 | NS | – | – | – |
BLOeM_4-115 | B2 III: | 9 | 26.92 | 4.45 | SB1 | 2.07 | NS | – | – | – |
BLOeM_4-116 | B1.5 III: | 9 | 33.62 | 5.76 | SB1 | 34.36 | NS | – | – | – |
BLOeM_5-002 | B2 IV | 9 | 184.80 | 10.02 | SB2 | 2.06 | S | y | 0.91773 | – |
BLOeM_5-003 | B2.5 III | 9 | 19.84 | 3.82 | RVvar | 10.28 | NS | – | – | – |
BLOeM_5-004 | B2.5 III-II | 9 | 12.76 | 3.09 | RVvar | 1.25 | NS | – | – | – |
BLOeM_5-011 | B2 III: | 9 | 22.74 | 3.84 | RVvar | 2.84 | NS | – | – | – |
BLOeM_5-013 | B2 III | 9 | 254.06 | 28.15 | SB2 | 1.56 | NS | y | 3.509807 | Misident. |
BLOeM_5-015 | B2 III: | 9 | 94.84 | 20.87 | SB1 | 500.00 | NS | – | – | – |
BLOeM_5-017 | B0.2 IV | 9 | 11.27 | 2.05 | RVvar | 1.40 | NS | – | – | – |
BLOeM_5-019 | B2 III | 9 | 44.77 | 20.97 | SB1 | 35.45 | NS | – | – | – |
BLOeM_5-024 | B0.7 III | 9 | 1.66 | 1.37 | RVcst | 16.34 | NS | – | – | – |
BLOeM_5-025 | B1 III | 5 | 8.19 | 1.72 | RVcst | 1.49 | NS | – | – | – |
BLOeM_5-027 | B1.5 III: | 9 | 167.13 | 25.61 | SB2 | 1.47 | S | – | – | – |
BLOeM_5-028 | B0 V | 9 | 27.40 | 3.53 | RVvar | 72.63 | NS | – | – | – |
BLOeM_5-030 | B1.5 III: | 9 | 39.21 | 5.84 | SB2 | 1.32 | NS | y | 2.638655 | Misident. - Half the OGLE P |
BLOeM_5-033 | B1 III | 9 | 4.97 | 1.54 | RVcst | 2.43 | NS | – | – | – |
BLOeM_5-037 | B0.5 V: | 9 | 53.73 | 12.87 | SB2 | 3.80 | NS | y | 2.722014 | Narrow cores - Misident. |
BLOeM_5-038 | B1.5 III-II | 9 | 19.79 | 11.81 | RVvar | 6.09 | NS | – | – | – |
BLOeM_5-039 | B1.5 III | 9 | 101.10 | 32.26 | SB1 | 33.35 | NS | – | – | – |
BLOeM_5-049 | B0.2 IV | 9 | 66.64 | 10.56 | SB1 | 4.98 | NS | y | 9.935441 | Half the OGLE P - SB2? |
BLOeM_5-051 | B0.2 III | 9 | 90.21 | 4.88 | SB1 | 9.49 | NS | – | – | – |
BLOeM_5-057 | B1 III: | 9 | 42.82 | 4.83 | SB2 | 14.21 | S | – | – | Possible SB3 - Similar to 3-086 |
BLOeM_5-058 | B1 III-II | 9 | 2.00 | 0.82 | RVcst | 1.32 | NS | – | – | – |
BLOeM_5-059 | B0.7 IV | 9 | 8.24 | 2.08 | RVvar | 8.70 | NS | – | – | – |
BLOeM_5-061 | B2 III: | 9 | 14.61 | 2.81 | RVvar | 500.00 | NS | – | – | – |
BLOeM_5-062 | B1.5 | 9 | 296.70 | 49.37 | SB3 | 79.96 | NS | y | 5.24094 | Affected by 3rd light |
BLOeM_5-063 | B0 V: | 4 | 264.01 | 57.31 | SB2 | 1.84 | NS | – | – | Too few epochs |
BLOeM_5-065 | B0 IV | 9 | 15.08 | 9.04 | RVvar | 500.00 | NS | – | – | – |
BLOeM_5-068 | B1 III-II | 9 | 17.27 | 8.18 | RVvar | 1.22 | NS | – | – | – |
BLOeM_5-070 | B2 III: | 9 | 9.10 | 2.41 | RVvar | 1.34 | NS | – | – | – |
BLOeM_5-074 | B0.7 IV | 9 | 22.11 | 8.43 | SB1 | 1.65 | S | – | – | – |
[…]